OpenStudy (anonymous):

Boolean algebra question: Does A'C'D' + A'C'D' stay as it is or does it merge into a single term? In other words, is the identity x + x = x applicable here?

OpenStudy (freckles):

x+x is 2x What is the meaning of A'C'D'?

OpenStudy (anonymous):

A' = not A. Boolean algebra.

OpenStudy (freckles):

what value is x,y,z are those suppose to be 1's and 0's

OpenStudy (anonymous):

x = 1 x' = 0

OpenStudy (freckles):

Well since you do have x+x you can use the identity 5 there

OpenStudy (freckles):

\[A'C'D'+A'C'D'=A'C'D'\] \[x+x=x\]

OpenStudy (anonymous):

x'y'z' =/= (xyz)' thought

OpenStudy (anonymous):

-t*

OpenStudy (anonymous):

x'y'z' = (x+y+z)'

OpenStudy (anonymous):

if it was (ACD)' + (ACD)' i wouldn't even be suspicious and write (ACD)' straight away

OpenStudy (freckles):

\[(A+C+D)'=A'C'D' \text{ assuming we can expand that the thingy using 15 } \\ \text{ but \in the \end this will still give us the same thing } \text{ \because }\\ (A+C+D)'+(A+C+D)'=(A+C+D)'= A'C'D'\]

OpenStudy (freckles):

but I kinda want to do some reasearch into this notation i'm not too faimilar with it and i'm assuming things I don't know anything about yet like i assume (A+C+D)'=A'C'D' I don't even know if (x+y)'=x'y' can be expanded to three variables

OpenStudy (freckles):

oh is + sign just the bit wise operation exclusive or?

OpenStudy (anonymous):

Yes it can be done even on whole expressions, it's called DeMorgan's law if you want to look it up.

OpenStudy (anonymous):

+ Is basically OR

OpenStudy (freckles):

now the only thing I need to know what is the meaning of xy

OpenStudy (freckles):

like is that and?

OpenStudy (freckles):

xy means x and y right

OpenStudy (anonymous):

Yes there's a "multiplication" between them, aka AND

OpenStudy (freckles):

so just for fun... A C D A' C' D' A' and C' and D' (A' and C' and D') or (A' and C' and D') T T T F F F F F T T F F F T F F T F F F T T F F T F T F T F F F F F T T T F F F F T F T F T F F F T T T F F F F F F F T T T T T As you see we have shown A'C'D'+A'C'D' =A'C'D' by truth table your initial thingy about applying x+x=x was correct

OpenStudy (anonymous):

Oh right! I know I was missing something, completely forgot about the truth table and k-map.

OpenStudy (anonymous):

Man thanks for responding I was really desperate with that one.

OpenStudy (freckles):

np I was just not used to the notation I was like omg what does this all mean

OpenStudy (anonymous):

That's some determination I must admit, diving in like that in a new subject.

OpenStudy (anonymous):

Just proved it's just A'C'D' using k-map as well.

OpenStudy (freckles):

well no I was just used to different notation